If a bulb of 60 W glows for 10 hours per day, what will be its energy consumption a day?
A. 1 KW
B. 6 KW
C. 0.6 KW

Grade:8

1 Answers

Arun
6 years ago
Dear Kushal
 
Energy consumption = 60* 10 * 1/1000 = 600/1000
 = 0.6 KWh
 
Regards
Arun (askIITians forum expert)